OpenStudy (anonymous):

what are the real solutions of the equation x^2-x+2=0

OpenStudy (anonymous):

\[ a=1\\b=-1\\c=2\\ \frac{-b\pm\sqrt{b^2-4ac}}{2a}\\ \frac{1\pm\sqrt{1-8}}{2}=\frac{1\pm\sqrt{-7}}{2} \]Because the square root of -7 is an imaginary number, this equation has no real solutions.

OpenStudy (anonymous):

polynomial is not factorable with real numbers